Adigirl29:

A company produces and sells solar panels for $520. The company's daily profit, P(x), can be modeled by the function P(x) = −6x2 + 156x + 1,000, where x is the number of $5 price increases for each solar panel. Use the graph to answer the questions. Part C: Identify the approximate value of the maximum of the function. Explain what the maximum of the function means in terms of the problem scenario.

Adigirl29:

1 attachment
Adigirl29:

this one is asking me for the value of the maximum of the function, I think it might be like 13 and 2000?

SoundwaveKun:

Here are the steps: Locate the vertex, calculate the maximum profit, and then solve for P(13). ^_^

Adigirl29:

okay so the vertex is 13,2000, for the profit how would I calculate it

SoundwaveKun:

@adigirl29 wrote:
okay so the vertex is 13,2000, for the profit how would I calculate it
Using this formula: Find the maximum value of the function by substituting x = 13 into the original function P(x):P(13) = -6(13^2) + 156(13) + 1,000 calculate P(13):P(13) = -6(169) + 2,028 + 1,000 P(13) = -1,014 + 2,028 + 1,000 P(13) = 2,028 + 1,000 P(13) = ?

Adigirl29:

I got 15028, is that right?

SoundwaveKun:

@adigirl29 wrote:
I got 15028, is that right?
Actually, the value should be $3,028.

Adigirl29:

oh got it

SoundwaveKun:

@adigirl29 wrote:
oh got it
Good job though! ^_^

Adigirl29:

tysm really this helped a lot

SoundwaveKun:

@adigirl29 wrote:
tysm really this helped a lot
Of course! 💙